Question 20 4 pts Find sum of the pair of complex numbers. 7-81, 4i 0 7-4i 07-121 11 - 81
Answer #1

Solve Giren Complex No. z = 7-8i, 4i then find sum of pair of Complex No. let z = 2 + Zz Zi= 7-si , 22=4i for sum of complex